Transform equation T=2pi√(L-h)/g into linear form as h=

1
Expert's answer
2021-11-07T19:28:34-0500

Expressing "h", find:


"\dfrac{L-h}{g} = \dfrac{T^2}{4\pi^2}\\
L-h = \dfrac{gT^2}{4\pi^2}\\
h = L - \dfrac{gT^2}{4\pi^2}"

Answer. "h = L - \dfrac{gT^2}{4\pi^2}"